Q1. 20 ml each of oxygen and hydrogen and 10ml of carbon monoxide are exploded in an enclosure. what will be the volume and compotion of the mixtures of the gases when they are cooled to room tempetature.
Q2. 224cm3 of ammonia undergoes catalytic oxidation in presence of Pt to give nitric oxide and water vapour. Calculate the volume of oxygen riquired for the reaction. All volumes measured at S.T.P
Q.3 Calculate the No. of Atom in 10 ltrs of ammonia (NH3) [ N=14,H=1].
1
Expert's answer
2013-02-15T05:54:20-0500
Q1.
For answering this you need to consider few reactions,which can go during explode:

2H2 + O2 = 2H2O

It means that all H2 reacts with half of O2

But you have another process here:

2CO + O2 = 2CO2

If you have ten ml of CO and ten ml of O2, the O2 is in excess.
Ten ml of CO reacts only with 5 ml of O2 with producing ten ml of CO2,so in the end you have 10 + 5 =15 ml of gases mixture.

Q2
& Reaction for this process is next.

4NH3 + 5O2 = 4NO + 6H2O

It means,that one volume of NH3 reacts with 5/4 volume of O2,so 224* 5/4=280 cm^3

Q3

& 10 ltrs N
NH3 -------4 atoms
22.4 6.02 *10^23


N=2.69 *10^23 atoms